OpenStudy (anonymous):

Solve 3x = –24. A. –27 B. –8 C. –12 D. –21 I got D is that correct?

OpenStudy (anonymous):

The whole point of these questions is to "isolate x". So what you do to one side you must do to the other. You want to cancel the 3 from the left by dividing it so it = 1, and if you divide on the left you must do it to the right.

OpenStudy (anonymous):

3x=-24 x=-24/3 = -8

OpenStudy (anonymous):

oh I thought I was supposed to subtract it from each side

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Thank you :) Can i ask one more question?

OpenStudy (anonymous):

For sure

OpenStudy (anonymous):

By what number must you multiply each side of the equation 3/4x=9 to produce the equivalent equation x = 12? A.-3/4 B.4/3 C. 4 D.1/4

OpenStudy (anonymous):

4/3 because you want to make the left side = 1.

OpenStudy (anonymous):